# A Theoretical Analysis of Detecting Large Model-Generated Time Series

### v2

@@ 2.1 Time Series Large Models
PatchTST (64words) introduces patchifying time series into embeddings, a technique later adopted in models such as GPT4TS (OneFitsAll) and Time-LLM (TimeLLM). These methods use patch embeddings as inputs to LLMs for prediction, but require extensive fine-tuning. Moreover, their performance and inference efficiency have been questioned (AreLLMUseful).
Recent work concentrates on models trained on vast and diverse time series datasets. MOMENT (moment) employs uniform random masking to generate patch embeddings for self-supervised pretraining. Chronos (Chronos) discretizes real-valued time series through scaling and quantization for forecasting and introduces data augmentation to improve generalization. Timer (timer) adopts a decoder-only transformer architecture, enabling multiple temporal tasks such as forecasting, anomaly detection, and imputation. Time-MoE (TimeMoE) leverages a Mixture-of-Experts framework to improve forecasting performance and enhance cross-domain generalization while maintaining computational efficiency. These methods collectively represent the emerging trend of general-purpose time series models capable of handling diverse tasks with minimal or no fine-tuning.
@@ 2.2 Model Generation Detection
The growing generative capabilities of LLMs have raised concerns about distinguishing model-generated content, particularly in the textual domain. Early work on model generation detection conducts supervised classification using bag of words (Release_Strategies) or neural representations (automatic; authorship), but often overfit and underperform on out-of-distribution data (deepfaketext). To overcome this limitation, zero-shot detectors use LLM output statistics, such as perplexity (relative_entropy) and log rank (gltr; unifying).
@@
Despite their success in text, these methods rely on the dense and discrete structure of language, making them unsuitable for continuous and information-sparse time series data. In this work, we extend zero-shot model generation detection to time series by introducing uncertainty-based metrics derived from model’s internal probability distributions.
@@ 3.1 Problem Formulation
We assume a white-box setting with access to the model’s internal probability distribution $p_{\theta}(\cdot|\mathbf{X}_{t})$ rather than the model architecture or parameters $\theta$. Our detection employs point-wise probabilistic TSLMs over an equidistant vocabulary $\mathcal{V}=\{v_{i}|v_{i+1}-v_{i}=\Delta,-R\leq v_{i}\leq R\}$ for new token sampling. Therefore, sampling token $v_{i}$ yields the numerical forecast $\hat{X}_{t+1}=v_{i}$. In the limit $\Delta\to 0$ and $R\to\infty$, $\mathcal{V}$ converges to $\mathbb{R}$ and induces a probability density $f_{\theta}$. Such a formulation allows for detection evaluation across diverse model architectures.
Some frequently used symbols throughout the paper are given in Table 1. In the following sections, we refer to distributions by their probability densities.
@@ 3.2 Preliminaries
$\mathrm{Eva}=\left[\sum_{i=1}^{\tau}g\left(\left|X_{i}-\hat{X}_{i}\right|^{p}\right)\right]^{1/p},1\leq p\leq\infty,$ (1)
@@ 4.2 Theoretical Analysis
Sampling-Induced Variance Scaling
@@
In particular, we derive a corollary from Section 4.2 that if metrics such as MSE are used as training loss, the contraction of uncertainty is immediate. {corollary} If the loss function is defined in the same form as the model evaluation function $\mathrm{Eva}$ in Eq. 1 (e.g., MSE),
we have $f_{\theta}(z=v)=\delta(v-T_{t})$.
@@
Recursive Variance Reduction
@@
Typically, the uncertainty scaling behavior is consistent and $\gamma_{t}<1$ or $\gamma_{t}>1$ for all $t$. When $\gamma_{t}<1$, the internal uncertainty, according to Eq. 2, is strictly lower than the weighted sum of historical uncertainties, leading to an exponential decay in uncertainty towards 0. By Section 4.2, the forecast series $\hat{X}_{t}$ converges to the trend $T_{t}$. Conversely, when $\gamma_{t}>1$, the uncertainty grows exponentially towards infinity, introducing excessive noise into the forecast that undermines trend capture and violates forecasting objectives. When $\gamma_{t}\equiv 1$, the uncertainty in forecast series is relatively stable for ideal models, which is unattainable for practical models due to noise accumulation.
@@ 5.1 Experimental Settings
Model and Datasets
We evaluate all detection methods using Chronos-T5 (large), a point-forecasting TSLM that discretizes continuous values over a fixed token vocabulary and produces probability distributions. We conduct experiments on 32 datasets across diverse domains (energy, finance, transportation), categorized as (1) in-distribution datasets, including 12 used during Chronos training; and (2) zero-shot datasets, consisting of 20 previously unseen datasets. The zero-shot setting evaluates the generalization of detection to unseen datasets to eliminate the effects of data leakage, and validates the “black-box” detection capability (see Section 5.3). For each dataset, we follow the default setup in Chronos to generate forecasts of horizon $H=64$, and use the corresponding $H$ observations as ground truth. We treat forecast series as positive (model-generated) samples and real series as negative samples.
Baselines
We compare UCE with 10 text-based baselines adapted for time series: (1) DNA‑GPT WScore (dnagpt), which measures the average log‑likelihood gap between regenerated and original text given a fixed prefix; (2) DetectGPT (detectgpt) and its efficient variant Fast‑DetectGPT (fast_detectgpt), which detect generation by measuring probability changes under model‑guided perturbations; (3) DetectLLM (detectllm), using Log‑Likelihood Log‑Rank Ratio (LRR) and Normalized Perturbed log‑Rank (NPR) metrics; (4) Intrinsic Dimension (Intrinsic_dimension), which classifies sequences by estimating the topological “intrinsic dimension” of token embeddings; (5) FourierGPT (FourierGPT), performing spectral analysis on token probability sequences; (6) Binocular (Binocular), leveraging two models to reduce the intrinsic perplexity of prefixes; and (7) Traditional metrics (gltr; automatic; Release_Strategies), including log‑likelihood $\log p(x)$, rank, and log‑rank. The details of the baselines are documented in the Appendix.
Evaluation Metrics
@@ 6.1 Idealized Model Assumption
To establish an architecture-agnostic detection methodology, we postulate a theoretically optimal Ideal Model. This abstraction avoids model-specific details to identify general behavior, but is mathematically unrealizable in practice. Existing TSLMs are approximations of this ideal, which undermine both the guarantee of perfect prediction (see Section 4.2) and the model’s ability to faithfully recover the true series distribution (see Section 4.2). Despite the limitation, practical models also exhibit similar uncertainty contraction behavior in forecasting.
@@ 6.2 Gaussian Noise Assumption
In this work, we assume Gaussian noise for analytical simplification: the sum of independent Gaussian random variables remains Gaussian, and the variance fully characterizes the distribution. Crucially, our proofs do not depend on the exact form of the Gaussian but only require the noise to be unimodal (greatest probability for trend), symmetric (equivalence for both positive and negative biases), and having a finite second moment (existence of variance). By the additivity of the variance for independent variables, our results naturally extend to such noise models. These properties are shared by most noise models commonly assumed in real-world time series analysis (e.g., Laplace), and the experimental results (see Appendix) validate the broad applicability of the contraction hypothesis under various noise types.
@@ 7 Conclusion
We investigate detecting TSLM-generated time series and hypothesize that they exhibit progressively decreasing uncertainty—unlike real data. Building on this, we propose the Uncertainty Contraction Estimator (UCE), which captures uncertainty to distinguish model-generated from real series and is validated both theoretically and empirically. Future work will further analyze the hypothesis under diverse model architectures and extend UCE to multivariate and batch-forecasting settings.

### v3

@@ 2.1 Time Series Large Models
PatchTST (64words) introduces splitting time series into patch embeddings, a technique later adopted in models such as GPT4TS (OneFitsAll) and Time-LLM (TimeLLM). These methods use patch embeddings as inputs to LLMs for prediction, but require extensive fine-tuning. Moreover, their performance and inference efficiency have been questioned (AreLLMUseful).
Recent work concentrates on models trained on vast and diverse time series datasets. MOMENT (moment) employs uniform random masking to generate patch embedding for self-supervised pretraining. Chronos (Chronos) discretizes real-valued time series through scaling and quantization for forecasting and introduces data augmentation to improve generalization. Timer (timer) adopts a decoder-only transformer architecture, enabling multiple temporal tasks such as forecasting, anomaly detection, and imputation. Time-MoE (TimeMoE) leverages a Mixture-of-Experts framework to improve forecasting performance and enhance cross-domain generalization while maintaining computational efficiency. These methods collectively represent the emerging trend of general-purpose time series models capable of handling diverse tasks with minimal or no fine-tuning.
@@ 2.2 Model Generation Detection
The growing generative capabilities of LLMs have raised concerns about distinguishing model-generated content, particularly in the textual domain. Early work on model generation detection conduct supervised classification using bag of words (Release_Strategies) or neural representations (automatic; authorship), but often overfit and underperform on out-of-distribution data (deepfaketext). To overcome this limitation, zero-shot detectors use LLM output statistics, such as perplexity (relative_entropy) and log rank (gltr; unifying).
@@
Despite their success in text, these methods rely on the dense and discrete structure of the language, making them unsuitable for continuous and information-sparse time series data. In this work, we extend zero-shot model generation detection to time series by introducing uncertainty-based metrics derived from model’s internal probability distributions.
@@ 3.1 Problem Formulation
We assume a white-box setting with access to the model’s internal probability distribution $p_{\theta}(\cdot|\mathbf{X}_{t})$ rather than to the model architecture or parameters $\theta$. Our detection employs point-wise probabilistic TSLMs over an equidistant vocabulary $\mathcal{V}=\{v_{i}|v_{i+1}-v_{i}=\Delta,-R\leq v_{i}\leq R\}$ for new token sampling. Therefore, sampling token $v_{i}$ yields the numerical forecast $\hat{X}_{t}=v_{i}$. In the limit $\Delta\to 0$ and $R\to\infty$, $\mathcal{V}$ converges to the real set $\mathbb{R}$ and a probability density function $f_{\theta}$ is obtained. Such a formulation allows for detection evaluation across diverse model architectures.
Some frequently used symbols throughout the paper are given in Table 1. In the following sections, we use probability density functions to refer to the distributions.
@@ 3.2 Preliminaries
$\mathrm{Eva}=\left[\sum_{i=1}^{\tau}g\left(\left|X_{i}-\hat{X}_{i}\right|^{p}\right)\right]^{\frac{1}{p}},1\leq p\leq\infty,$ (1)
@@ 4.2 Theoretical Analysis
Sampling-Induced Variance Scaling.
@@
In particular, we derive a corollary from Section 4.2 that if metrics such as MSE are used as training loss, the contraction of uncertainty is immediate. {corollary} If the loss function is defined in the same form as the model evaluation function $\mathrm{Eva}$ in Eq. 1 (e.g., MSE), we have $f_{\theta}(z=v)=\delta(v-T_{t})$.
@@
Recursive Variance Reduction.
@@
Typically, the uncertainty scaling behavior is consistent and $\gamma_{t}<1$ or $\gamma_{t}>1$ for all $t$. When $\gamma_{t}<1$, the internal uncertainty, according to Eq. 2, is strictly lower than the weight sum of historical uncertainties, leading to an exponential decay in uncertainty towards 0. By Section 4.2, the forecast series $\hat{X}_{t}$ converges to the trend $T_{t}$. Conversely, when $\gamma_{t}>1$, the uncertainty grows exponentially towards infinity, introducing excessive noise into the forecast that undermines trend capture and violates forecasting objectives. When $\gamma_{t}\equiv 1$, the uncertainty in forecast series is relatively stable for ideal models, which is unattainable for practical models due to noise accumulation.
@@ 5.1 Experimental Settings
Model and Datasets.
We evaluate all detection methods using Chronos-T5 (large), a point forecasting TSLM that discretizes continuous values over a fixed token vocabulary and produces probability distributions. We conduct experiments on 32 datasets across diverse domains (energy, finance, transportation), categorized as (1) in-distribution datasets, including 12 used during Chronos training; and (2) zero-shot datasets, consisting of 20 previously unseen datasets. The zero-shot setting evaluates the generalization of detection to unseen datasets to eliminate the effects of data leakage, and validates the “black-box” detection capability (see Section 5.3). For each dataset, we follow the default setup in Chronos to generate forecasts of horizon $H=64$, and use the corresponding $H$ observations as ground truth. We treat forecast series as positive (model-generated) samples and real series as negative samples.
Baselines.
We compare UCE with 10 text-based baselines adapted for time series: (1) DNA‑GPT WScore (dnagpt), which measures the average log‑likelihood gap between regenerated and original text given a fixed prefix; (2) DetectGPT (detectgpt) and its efficient variant Fast‑DetectGPT (fast_detectgpt), which detect generation by measuring probability changes under model‑guided perturbations; (3) DetectLLM (detectllm), using Log‑Likelihood Log‑Rank Ratio (LRR) and Normalized Perturbed log‑Rank (NPR) metrics; (4) Intrinsic Dimension (Intrinsic_dimension), which classifies sequences by estimating the topological “intrinsic dimension” of token embeddings; (5) FourierGPT (FourierGPT), performing spectral analysis on token probability sequence; (6) Binocular (Binocular), leveraging two models to reduce the intrinsic perplexity of prefixes; and (7) Traditional metrics (gltr; automatic; Release_Strategies), including log‑likelihood $\log p(x)$, rank, and log‑rank. The details of the baselines are documented in the Appendix.
Evaluation Metrics.
@@ 6.1 Idealized Model Assumption
To establish an architecture-agnostic detection methodology, we postulate a theoretically optimal Ideal Model. This abstraction avoids model-specific details to identify general behaviors, but is mathematically unrealizable in practice. Existing TSLMs are approximations of this ideal, which undermine both the guarantee of perfect prediction (see Section 4.2) and the model’s ability to faithfully recover the true series distribution (see Section 4.2). Despite the limitation, practical models also exhibit similar uncertainty contraction behaviors in forecasting.
@@ 6.2 Gaussian Noise Assumption
In this work, we assume Gaussian noise for analytical simplification: the sum of independent Gaussian noises remains Gaussian, and the variance fully characterizes the distribution. Crucially, our proofs do not depend on the exact form of the Gaussian but only require the noise to be unimodal (greatest probability for trend), symmetric (equivalence for both positive and negative biases), and having a finite second moment (existence of variance). By the additivity of the variance for independent variables, our results naturally extend to such noise models. These properties are shared by most noise models commonly assumed in real-world time series analysis (e.g., Laplacian), and the experimental results (see Appendix) validate the broad applicability of the contraction hypothesis under various noise types.
@@ 7 Conclusion
We investigate detecting TSLM‑generated time series and hypothesize that they exhibit progressively decreasing uncertainty–unlike real data. Building on this, we propose the Uncertainty Contraction Estimator (UCE), which captures uncertainty to distinguish model-generated from real series and is validated both theoretically and empirically. Future work will further analyze the hypothesis under diverse model architectures and extend UCE to multivariate and batch-forecasting settings.